Solve the following equations and check your results  x = 4/5 (x + 10)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x=4/5(x+10)
⇒x=4x/5+40/5
⇒x–(4x/5)=8
⇒(5x–4x)/5=8
⇒x=8×5
⇒x=40
Putting the value of x in RHS and LHS we get,
40=4/5(40+10)
⇒40=4/5×50
⇒40=200/5
⇒40=40
⇒LHS=RHS

Thanks